# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ Appendix E The geometry and complexity of Baumslag-Solitar groups
We now turn to examining the average-case complexity of words in ${\sf BS}$ (or more precisely, the complexity of typical words in ${\sf BS}$). Unfortunately, only a few results on average-case Dehn functions are known (one example for nilpotent groups is Ref. Young (2008)), and average-case complexity has not been studied for groups with superpolynomial Dehn function. We emphasize that our analysis is heuristic, relying on several reasonable claims backed up by numerics in some places. Letting $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu(L)$ denote the Dehn function of a typical word in $K_{e,L}$, we claim
@@
Proving this claim—which we believe is within reach—constitutes an interesting direction for future research.
@@
Proposition 16 (Burillo and Elder (2015)).
Any word can always be mapped to a unique standard form
@@
  • If $B\not\in\mathbb{Z}$, then $k>0$, and $k$ is determined by the number of significant figures after the decimal point when $B$ is represented in binary, [^37] There is a small subtlety here, since this approach is incorrect if $l=0$ and $n=2^{m}n_{o}$ with $m>0,n_{o}\in 2\mathbb{Z}+1$. However in this case it is easy to show that the naive values of $(k,n,l)$ one obtains from this procedure in this case are $(k,n,l)_{\rm naive}=(k-m,n_{0},-m)$. Since $l_{\rm naive}<0$, this mistake is easy to identify and correct for. after which both $l$ and $n$ are determined.
@@
Numerically implementing the above procedure for several values of $L$ gives the histogram shown in Fig. 12, where for simplicity we plot an estimate of $|g(w)|$ as $k+l+\log_{2}(|n|+1)$, which has the same asymptotic scaling as the true value of the geodesic. From this we see that the probability of getting $|g(w)|=0$ (viz. getting $w\in K_{e}$) is suppressed (we have already seen that ${\rm Pr}(w\in K_{e})\sim e^{-L^{1/3}}$), and that the typical geodesic distance goes as $|g(w)|\sim\sqrt{L}$.
@@
Since we are sampling random words, $n_{b}(w)=k-l$ will converge to a Gaussian of width $\sqrt{L}$; thus a typical word will reach a depth of $\sqrt{L}$ on the $b$-tree part of the Cayley graph. We also expect the $k+l$ contribution to the geodesic estimate (163) to scale as $\sqrt{L}$; indeed this can be numerically verified to be the case. The fact that a random word typically has $|g(w)|\sim\sqrt{L}$ then means that the $\log|n|$ contribution to the geodesic estimate scales as $\log|n|\lesssim\sqrt{L}$. This means that words which traverse exponentially far along the $a$ axis of the Cayley graph (viz. those with $\log|n|\sim L$) are rare, meaning that we should not expect the Dehn time of typical words to be close to the worst-case result. To determine whether $\log|n|<\sqrt{L}$ or $\log|n|\sim\sqrt{L}$, we simply make a histogram of $(\log|n|)/\sqrt{L}$, with the result shown in Fig. 13. We see that the $\log|n|$ part makes an $O(1)$ contribution to the expected geodesic distance, and we learn that
@@
Before moving on, we note as an aside that a randomly chosen word in $K_{e,L}$ is numerically observed to be essentially homogeneous as far as the conserved $n_{\mathtt{b}}$ density is concerned, meaning that $\frac{1}{|K_{e,L}|}\sum_{w\in K_{e,L}}\langle w|n_{\mathtt{b},i}|w\rangle\approx 0$ for all $i$ (although as far as we can tell there is no symmetry which enforces this expectation value to vanish identically). This is demonstrated in Fig. 14, and is important for numerically determining thermalization times in the manner of Sec. IV.5.
@@
The rough intuition leading to claim 1 is as follows. Let $L\in 2\mathbb{N}$ for simplicity and consider a random word $w=w_{L}w_{R}$ in $K_{e,L}$, with $|w_{L}|=|w_{R}|=L/2$. Then to the extent that $w_{L,R}$ behave like random length-$L/2$ words, the midpoint of the loop defined by $w$ will be at a distance $\sim 2^{\sqrt{L}}$ along the $a$ axis from the origin, as follows from observation 1. Given this, we may expect that area of $w$ also scales as $2^{\sqrt{L}}$. The only exception is if $w_{R}$ is very close to $w_{L}^{-1}$ which implies that $w$ subtends little area, but we will argue that such an event is unlikely. We begin with an argument about certain types of words which we will argue are likely to occur as subwords of typical elements in $K_{e,L}$:
@@
Consider a length-$L$ word $w$ chosen randomly subject to the following constraints. First, $n_{b}(w)=0$. Second, the returning walk induced on $\mathbb{Z}$ by restricting to the $b,b^{-1}$ characters of $w$ is constrained to $\mathbb{Z}^{\geq 0}$ (i.e. it is a Dyck walk), meaning that the cumulative sums $n_{b}(x)=\sum_{j=1}^{x}(\delta_{w_{i},b}-\delta_{[w_{i}]_{j},b})$ are positive for all $x$. Let $S_{Dyck}(w)$ be the set of all length-$L$ words $w^{\prime}\sim w$ obeying the above constraint. For $w^{\prime}$ is drawn randomly from $S_{Dyck}(w)$, $w$ and $w^{\prime}$ satisfy $d(w,w^{\prime})\sim 2^{\sqrt{L}}$ with high probability.
Argument: The Dyck walk property means that all $w^{\prime}\in S_{Dyck}(w)$ are reducible to $a^{n_{w}}$ for some $n_{w}$, with $n_{w}$ shared by all words in $S_{Dyck}(w)$. This can be seen by an inspection of the Cayley graph, or by recalling the canonical form (162) (the walks corresponding to a group element are read right to left, so $b^{k}a^{n}b^{-l}\in S_{Dyck}(w)$ only if $l=k=0$). Since $w$ was chosen randomly from the set of words obeying the Dyck walk condition, the $b$-walk defined by $w$ is exponentially likely to reach a height of $O(\sqrt{L})$, viz. to have $\max_{x}n_{b}(x)\sim\sqrt{L}$. A random $w$ fullfiling this condition can be seen to be exponentially likely to reduce to a word $a^{n_{w}}$ of length $n_{w}\sim 2^{\sqrt{L}}$, as can be seen by e.g. following the procedure that brings $w$ into canonical form (162).
@@
Supose now that $n_{w}$ scales as $2^{\sqrt{L}}$, and consider a random element $w^{\prime}\in S_{Dyck}(w)$. We claim that the distance between $w$ and $w^{\prime}$ is exponentially likely to scale as $d(w,w^{\prime})\sim 2^{\sqrt{L}}$. This is because $w,w^{\prime}$ are exponentially likely to travel on different sheets of the Cayley graph for nearly all of the length of their walks. More precisely, consider the projection of $w,w^{\prime}$ onto the b-tree, and define the branch point $Br(w,w^{\prime})$ as the largest depth of a vertex in the tree visited by both $w$ and $w^{\prime}$. We claim that $Br(w,w^{\prime})$ is exponentially likely to be $O(1)$, so that $w,w^{\prime}$ indeed travel on separate sheets for nearly the entirety of their trajectories. To demonstrate this more carefully, consider how returning random walks on ${\mathsf{CG}}_{\sf BS}$ behave when projected onto the b-tree. When the walk on the tree moves to larger scales of ${\mathsf{CG}}_{\sf BS}$, we will refer to it as moving “upwards”, and when it moves to smaller scales we will say that it moves “downwards”. At a given vertex in ${\mathsf{CG}}_{\sf BS}$, moving upwards while staying on the same sheet can be done by moving directly upwards, or by moving to the left or right by an even number of steps, and then moving up. Moving upwards onto a different sheet on the other hand is done by moving left or right by an odd number of steps before moving up. Finally, moving downwards (on the same sheet) can be done by moving an arbitrary amount either left or right, and then moving downwards. This means that the probability to move downwards is equal to the probability of moving upwards, despite the fact that moving upwards can be done on either of two sheets. More precisely, let $p^{\nwarrow},p^{\nearrow}$ and $p^{\downarrow}$ be the probabilities of moving up on the same sheet, up on a different sheet, and down, respectively. Then
$p^{\nwarrow}$ $=\frac{4}{5}\frac{1}{4}\sum_{k\in\mathbb{Z}}\sum_{l=|k|}^{\infty}{2l\choose l+|k|}(1/4)^{2l}=\frac{1}{3}$ (166)
@@
The fact that $p^{\downarrow}=1/2$ means that as far as motion in the tree is concerned, the walk will not move ballistically upwards or downwards, but will instead move diffusively. Using the above transition probabilities, the expected behavior of $Br(w,w^{\prime})$ can be calculated analytically using generating functions. The details are rather messy however, and since we regard the claim of $Br(w,w^{\prime})$ as being intuitively rather clear, we will be content with a numerical demonstration of this result, which we provide in Fig. 15. Now we return to our discussion of the distance between $w$ and $w^{\prime}$. As just argued, $Br(w,w^{\prime})$ is exponentially likely to be $O(1)$. The contractibility of ${\mathsf{CG}}_{\sf BS}$ means that the minimal bounding surface linking $w$ to $w^{\prime}$ must consist of all cells bounded by $w$ and the $a$ axis that lie at a depth greater than $Br(w,w^{\prime})$, together with the analogous set of cells for $w^{\prime}$ (a similar argument arose in the proof of proposition 13, where there we effectively had $Br(w,w^{\prime})=0$). Since each of these contributions to the bounding surface consists of $\sim 2^{\sqrt{L}}$ cells, we indeed have $d(w,w^{\prime})\sim 2^{\sqrt{L}}$. ∎ We are now in a position to argue for the correctness of claim 1. While Dyck walks do not constitute a constant fraction of all returning walks [^38] The number of length-$L$ Dyck walks scales as $\sim L^{-3/2}2^{L}$, while the number of returning walks goes as $\sim L^{-1/2}2^{L}$; thus the fraction of returning walks that are also Dyck walks scales as $\sim 1/L$., the basic idea is to realize that a generic word in $K_{e}$ is likely to contain at least two subwords of size $\sim\sqrt{L}$ obeying the Dyck property, allowing an application of the above argument. First, when sampling a random word in $K_{e}$, we can first sample uniformly from all $5^{L/2}$ words $w_{L}$ of length $L/2$, and then sample from all words $w_{R}$ of length $L/2$ such that $w_{L}w_{R}\sim e$. Since $w_{L}$ is chosen randomly, we expect that with unit probaility in the $L\rightarrow\infty$ limit, the walk defined by $w_{L}$ reaches a distance of order $\sim 2^{\sqrt{L}}$ along the $a$ axis of the Cayley graph (this was numerically demonstrated to be the case in Fig. 13). Since $|w_{L}|\sim L$, reaching this distance is only possible if $w_{L}$ contains an excursion along a particular sheet of the $b$-tree which reaches a maximal depth of $\sim\sqrt{L}$. $w_{L}$ must therefore contain at least one subword $w_{L,D}$ which performs a Dyck walk of height $\sim\sqrt{L}$. Given $w_{L}$, we now consider sampling words $w_{R}$ such that $w_{L}w_{R}\sim e$. Since $w_{R}$ must also traverse a distance of $\sim 2^{\sqrt{L}}$ along the $a$ axis, it too must contain a subword $w_{R,D}$ which performs a Dyck walk of height $\sim\sqrt{L}$. By the contractibility of the Cayley 2-complex, the areas that $w_{L,D}$ and $w_{R,D}$ define with respect to the $a$ axis can cancel out only if the excursions that $w_{L,D}$ and $w_{R,D}$ perform occur along the same sheets of the $b$-tree for a fraction of their respective walks exponentially close to 1. The chance for this to occur is however exponentially small in the depth of the walk (since the number of such walks on the b-tree grows exponentially in their length), which goes as $\sqrt{L}$. Therefore the areas contributed by $w_{L,D}$ and $w_{R,D}$ are exponentially unlikely to cancel, and thus the area of $w=w_{L}w_{R}$ will scale as $2^{\sqrt{L}}$ with probability approaching 1 as $L\rightarrow\infty$.
@@
We end our argument for claim 1 by providing a degree of evidence from numerics which supports the conclusions of the above arguments. We do this by considering the distance that randomly chosen words in $K_{e,L}$ travel from the origin. More precisely, for each randomly-chosen $w\in K_{e,L}$, we subdivide $w$ into two equal-length halves as $w=w_{L}w_{R}$ and then compute the typical amount that $w_{L}$ extends along the $a$ axis of ${\mathsf{CG}}_{\sf BS}$. Sampling uniformly from $K_{e,L}$ is done by sampling random words and then post-selecting on them being in $K_{e,L}$, which for large $L$ is numerically costly on account of the postselection succeeding with probably $\sim e^{-\alpha L^{1/3}}$. This limits our numerics to relatively modest system sizes, for which finite size effects are rather strong. Nevertheless, the results shown in Fig. 16 shown that typical words in $K_{e,L}$ have $w_{L}$ which indeed extends a distance of $\sim 2^{\sqrt{L}}$ along the $a$ axis of ${\mathsf{CG}}_{\sf BS}$. As argued above, this implies the correctness of claim 1 (as the only way for this to be false is if $w_{L}$, $w_{R}$ almost always traverse the same path of the b-tree, which we have already demonstrated is exponentially unlikely to be the case). ∎ Finally, we note that by following similar logic as in the proof of proposition 4, the same scaling as in claim 1 can be argued to occur for randomly chosen words, and not just those in $K_{e}$. More precisely, from claim 1 we expect that if $w$ is a random length-$L$ word and $w^{\prime}$ is subsequently randomly chosen from $K_{g(w)}(L)$, then $d(w,w^{\prime})\sim 2^{\sqrt{L}}$ with constant probability.

### 11 Apr 2024 v2

@@ Appendix E The geometry and complexity of Baumslag-Solitar groups
We now turn to examining the average-case complexity of words in ${\sf BS}$ (or more precisely, the complexity of typical words in ${\sf BS}$). Unfortunately, only a few results on average-case Dehn functions are known (one example for nilpotent groups is Ref. [45]), and average-case complexity has not been studied for groups with superpolynomial Dehn function. We emphasize that our analysis in this section is not rigorous and relies on several reasonable claims backed up by some numerical evidence. Letting $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu(L)$ denote the Dehn function of a typical word in $K_{e,L}$, we claim
@@
Rigorously proving this claim—which we believe is within reach—constitutes an interesting direction for future research.
@@
Proposition 16 ([88]).
Any word in ${\sf BS}$ can always be mapped to a unique standard form
@@
  • If $B\not\in\mathbb{Z}$, then $k>0$, and $k$ is determined by the number of significant figures after the decimal point when $B$ is represented in binary, [^38] There is a small subtlety here, since this approach is incorrect if $l=0$ and $n=2^{m}n_{o}$ with $m>0,n_{o}\in 2\mathbb{Z}+1$. However in this case it is easy to show that the naive values of $(k,n,l)$ one obtains from this procedure in this case are $(k,n,l)_{\rm naive}=(k-m,n_{0},-m)$. Since $l_{\rm naive}<0$, this mistake is easy to identify and correct for. which then allows one to determine $l$ and $n$.
@@
Numerically implementing the above procedure for several values of $L$ gives the histogram shown in Fig. 13, where for simplicity we plot an estimate of $|g(w)|$ as $k+l+\log_{2}(|n|+1)$, which has the same asymptotic scaling as the true value of the geodesic. From this we see that the probability of getting $|g(w)|=0$ (namely getting $w\in K_{e}$) is suppressed (we have already seen that ${\rm Pr}(w\in K_{e})\sim e^{-L^{1/3}}$), and that the typical geodesic distance goes as $|g(w)|\sim\sqrt{L}$.
@@
Since we are sampling random words, $n_{b}(w)=k-l$ will converge to a Gaussian of width $\sqrt{L}$; thus a typical word will reach a depth of $\sqrt{L}$ on the $b$-tree part of the Cayley graph. We also expect the $k+l$ contribution to the geodesic estimate (168) to scale as $\sqrt{L}$; indeed this can be numerically verified to be the case. The fact that a random word typically has $|g(w)|\sim\sqrt{L}$ then means that the $\log|n|$ contribution to the geodesic estimate scales as $\log|n|\lesssim\sqrt{L}$. This means that words which traverse exponentially far along the $a$ axis of the Cayley graph (namely those with $\log|n|\sim L$) are rare, meaning that we should not expect the Dehn time of typical words to be close to the worst-case result. To determine whether $\log|n|<\sqrt{L}$ or $\log|n|\sim\sqrt{L}$, we simply make a histogram of $(\log|n|)/\sqrt{L}$, with the result shown in Fig. 14. We see that the $\log|n|$ part adds at most an $O(1)$ contribution to the expected geodesic distance, resulting in:
@@
Before moving on, we note as an aside that a randomly chosen word in $K_{e,L}$ is numerically observed to be essentially homogeneous as far as the conserved $n_{\mathtt{b}}$ density is concerned, meaning that $\frac{1}{|K_{e,L}|}\sum_{w\in K_{e,L}}\langle w|n_{\mathtt{b},i}|w\rangle\approx 0$ for all $i$ (although as far as we can tell there is no symmetry which enforces this expectation value to vanish, so we do not have a clear explanation for this). This is demonstrated in Fig. 15, and is important for numerically determining thermalization times in the manner of Sec. IV.5.
@@
The rough intuition leading to claim 1 is as follows. Let $L\in 2\mathbb{N}$ for simplicity and consider a random word $w=w_{L}w_{R}$ in $K_{e,L}$, with $|w_{L}|=|w_{R}|=L/2$. Then to the extent that $w_{L,R}$ behave like random length-$L/2$ words, the midpoint of the loop defined by $w$ will be at a distance $\sim 2^{\sqrt{L}}$ along the $a$ axis from the origin, as follows from observation 1. Given this, we may expect that area of $w$ also scales as $2^{\sqrt{L}}$. The only exception is if $w_{R}$ follows a path which is close to $w_{L}^{-1}$ which implies that $w$ subtends little area, but we will argue that such an event is unlikely. Intuitively, this is because the probability the $w_{R}$ and $w_{L}^{-1}$ follow different branches should be large. We begin with an argument about certain types of words which we will argue are likely to occur as subwords of typical elements in $K_{e,L}$:
@@
Consider a length-$L$ word $w$ chosen randomly subject to the following constraints. First, $n_{b}(w)=0$. Second, the returning walk induced on $\mathbb{Z}$ by restricting to the $b,b^{-1}$ characters of $w$ is constrained to $\mathbb{Z}^{\geq 0}$ (i.e. it is a Dyck walk). Put another way, the cumulative sums $n_{b}(x)=\sum_{j=1}^{x}(\delta_{w_{i},b}-\delta_{[w_{i}]_{j},b})$ are positive for all $x$. Let $S_{Dyck}(w)$ be the set of all length-$L$ words $w^{\prime}\sim w$ obeying the above constraint. For $w^{\prime}$ is drawn uniformly from $S_{Dyck}(w)$, $w$ and $w^{\prime}$ satisfy $d(w,w^{\prime})\sim 2^{\sqrt{L}}$ with high probability.
Argument: The Dyck walk property means that all $w^{\prime}\in S_{Dyck}(w)$ are reducible to $a^{n_{w}}$ for some $n_{w}$, with $n_{w}$ the same for all words in $S_{Dyck}(w)$. This can be seen by an inspection of the Cayley graph, or by recalling the canonical form (167) (the walks corresponding to a group element are read right to left, so $b^{k}a^{n}b^{-l}\in S_{Dyck}(w)$ only if $l=k=0$). Since $w$ was chosen randomly from the set of words obeying the Dyck walk condition, the $b$-walk defined by $w$ is (using simple concentration inequalities) likely to reach a height of $O(\sqrt{L})$, namely to have $\max_{x}n_{b}(x)\sim\sqrt{L}$. A random $w$ fulfilling this condition can be seen to be likely to reduce to a word $a^{n_{w}}$ of length $n_{w}\sim 2^{\sqrt{L}}$, which can be seen by following the procedure bringing $w$ into the canonical form (167).
@@
Supose now that $n_{w}$ scales as $2^{\sqrt{L}}$, and consider a random element $w^{\prime}\in S_{Dyck}(w)$. We claim that the distance between $w$ and $w^{\prime}$ is likely to scale as $d(w,w^{\prime})\sim 2^{\sqrt{L}}$. This is because $w,w^{\prime}$ are likely to travel on different sheets of the Cayley graph for nearly all of the length of their walks. More precisely, consider the projection of $w,w^{\prime}$ onto the b-tree, and define the branch point $Br(w,w^{\prime})$ as the largest depth of a vertex in the tree shared by the paths undertaken by both $w$ and $w^{\prime}$. We claim that $Br(w,w^{\prime})$ is exponentially likely to be $O(1)$, so that $w,w^{\prime}$ indeed travel on separate sheets for nearly the entirety of their trajectories. To demonstrate this more carefully, consider how returning random walks on ${\mathsf{CG}}_{\sf BS}$ behave when projected onto the b-tree. When the walk on the tree moves to larger scales of ${\mathsf{CG}}_{\sf BS}$, we will refer to it as moving “upwards”, and when it moves to smaller scales we will say that it moves “downwards”. At a given vertex in ${\mathsf{CG}}_{\sf BS}$, moving upwards while staying on the same sheet can be done by moving directly upwards, or by moving to the left or right by an even number of steps, and then moving up. Moving upwards onto a different sheet on the other hand is done by moving left or right by an odd number of steps before moving up. Finally, moving downwards (on the same sheet) can be done by moving an arbitrary amount either left or right, and then moving downwards. The above reasoning shows that the probability to move downwards is equal to the probability of moving upwards, despite the fact that moving upwards can be done on either of two sheets. More precisely, let $p^{\nwarrow},p^{\nearrow}$ and $p^{\downarrow}$ be the probabilities of moving up on the same sheet, up on a different sheet, and down, respectively. Then
$p^{\nwarrow}$ $=\frac{4}{5}\frac{1}{4}\sum_{k\in\mathbb{Z}}\sum_{l=|k|}^{\infty}{2l\choose l+|k|}\left(\frac{1}{4}\right)^{2l}=\frac{1}{3}$ (171)
@@
The fact that $p^{\downarrow}=1/2$ means that as far as motion in the tree is concerned, the walk will not move ballistically upwards or downwards, but will instead move diffusively. Using the above transition probabilities, the expected behavior of $Br(w,w^{\prime})$ can be calculated analytically using generating functions. The details are rather messy however, and we content ourselves with a numerical demonstration that $Br(w,w^{\prime})=O(1)$ with constant probability, see Fig. 16. Now we return to our discussion of the distance between $w$ and $w^{\prime}$. As just argued, $Br(w,w^{\prime})$ is likely to be $O(1)$. The contractibility of ${\mathsf{CG}}_{\sf BS}$ means that the minimal bounding surface linking $w$ to $w^{\prime}$ must consist of all cells bounded by $w$ and the $a$ axis that lie at a depth greater than $Br(w,w^{\prime})$, together with the analogous set of cells for $w^{\prime}$ (a similar argument arose in the proof of proposition 13, whereby in that case $Br(w,w^{\prime})=0$). Since each of these contributions to the bounding surface consists of $\sim 2^{\sqrt{L}}$ cells, we thus have $d(w,w^{\prime})\sim 2^{\sqrt{L}}$. ∎ We are now in a position to argue for the correctness of Claim 1. Dyck walks do not constitute a constant fraction of all returning walks: the number of length-$L$ Dyck walks scales as $\sim L^{-3/2}2^{L}$), while the number of returning walks goes as $\sim L^{-1/2}2^{L}$. Thus only a fraction $\sim 1/L$ of returning walks are also Dyck walks. Nevertheless, this still gives us enough information to argue that the expected Dehn function is $\gtrsim 2^{\sqrt{L}}$. To show that the Dehn function is $\sim 2^{\sqrt{L}}$ with constant probability, the basic idea is to realize that a generic word in $K_{e}$ is likely to contain at least two subwords of size $\sim\sqrt{L}$ obeying the Dyck property, allowing for application of the above argument. First, when sampling a random word in $K_{e}$, we can first sample uniformly from all $5^{L/2}$ words $w_{L}$ of length $L/2$, and then sample from all words $w_{R}$ of length $L/2$ such that $w_{L}w_{R}\sim e$. Since $w_{L}$ is chosen randomly, we expect that with unit probability in the $L\rightarrow\infty$ limit, the walk defined by $w_{L}$ reaches a distance of order $\sim 2^{\sqrt{L}}$ along the $a$ axis of the Cayley graph (this was numerically demonstrated to be the case in Fig. 14). Since $|w_{L}|\sim L$, reaching this distance is only possible if $w_{L}$ contains an excursion along a particular sheet of the $b$-tree which reaches a maximal depth of $\sim\sqrt{L}$. $w_{L}$ must therefore contain at least one subword $w_{L,D}$ which performs a Dyck walk of height $\sim\sqrt{L}$. Given $w_{L}$, we now consider sampling words $w_{R}$ such that $w_{L}w_{R}\sim e$. Since $w_{R}$ must also traverse a distance of $\sim 2^{\sqrt{L}}$ along the $a$ axis, it too must contain a subword $w_{R,D}$ which performs a Dyck walk of height $\sim\sqrt{L}$. By the contractibility of the Cayley 2-complex, the areas that $w_{L,D}$ and $w_{R,D}$ define with respect to the $a$ axis can cancel out only if the excursions that $w_{L,D}$ and $w_{R,D}$ perform occur along the same sheets of the $b$-tree for a fraction of their respective walks close to 1. The chance for this to occur is however exponentially small in the depth of the walk (since the number of such walks on the b-tree grows exponentially in their length), which goes as $\sqrt{L}$. Therefore the areas contributed by $w_{L,D}$ and $w_{R,D}$ are unlikely to cancel, and thus the area of $w=w_{L}w_{R}$ will scale as $2^{\sqrt{L}}$ with probability approaching 1 as $L\rightarrow\infty$.
@@
We end our argument for claim 1 by providing a degree of evidence from numerics which supports the conclusions of the above arguments. We do this by considering the distance that randomly chosen words in $K_{e,L}$ travel from the origin. More precisely, for each randomly-chosen $w\in K_{e,L}$, we subdivide $w$ into two equal-length halves as $w=w_{L}w_{R}$ and then compute the typical amount that $w_{L}$ extends along the $a$ axis of ${\mathsf{CG}}_{\sf BS}$. Sampling uniformly from $K_{e,L}$ is done by sampling random words and then postselecting on them being in $K_{e,L}$, which for large $L$ is numerically costly on account of the postselection succeeding with probably $\sim e^{-\alpha L^{1/3}}$. This limits our numerics to relatively modest system sizes, for which finite size effects are rather strong. Nevertheless, the results shown in Fig. 17 shown that typical words in $K_{e,L}$ have $w_{L}$ which indeed extends a distance of $\sim 2^{\sqrt{L}}$ along the $a$ axis of ${\mathsf{CG}}_{\sf BS}$. As argued above, this implies the correctness of claim 1 (since this is false only if $w_{L}$, $w_{R}$ almost always traverse the same path of the b-tree, which we have already demonstrated is unlikely). ∎ Finally, we note that by following similar logic as in the proof of proposition 4, the same scaling as in claim 1 can be argued to occur for randomly chosen words, and not just those in $K_{e}$. More precisely, from claim 1 we expect that if $w$ is a random length-$L$ word and $w^{\prime}$ is subsequently randomly chosen from $K_{g(w)}(L)$, then $d(w,w^{\prime})\sim 2^{\sqrt{L}}$ with constant probability.